OpenStudy (aryavegapunk):

f(x)= c^2 +7x if x<4. f(x)= x^3-cx if x>/=4. What is C?

OpenStudy (aryavegapunk):

\[f(x)= cx ^{2}+7x \rightarrow x<4\] \[f(x)=x ^{3}-Cx \rightarrow x \ge 4\]

OpenStudy (anonymous):

plug in the 4, then set the 2 pieces of the function equal to solve for c

OpenStudy (anonymous):

\[16c+28=64-4c\]

OpenStudy (aryavegapunk):

Continuous on \[\left( -\infty, \infty\right)\] ?

OpenStudy (anonymous):

that's how you would make f(x) continuous, by making the 2 pieces equal at 4

OpenStudy (aryavegapunk):

Thank you!!

OpenStudy (anonymous):

you're welcome